OpenStudy (ledah):

Solve for n: n + (-5) = 12

Parth (parthkohli):

Aha \( \color{Black}{\Rightarrow n - 5 = 12}\)

Parth (parthkohli):

Add 5 to both sides, I'm afraid that's the stage I am going to leave you to do it on your own :P

OpenStudy (anonymous):

n=12+5 n=17

OpenStudy (anonymous):

got it?

OpenStudy (ledah):

understood. one minute

OpenStudy (anonymous):

ya speak up?